OpenStudy (anonymous):

mathew loaned three friends a total of 27 dollars. All he can remember is that he loaned the second friend $5 more than the first and the third friend twice as much as the second. Write and solve an equation to determined how much he loaned each friend

OpenStudy (anonymous):

Let the amt loaned to the second be s first be f third be t we know s = t+5 t=2s f+s+t = 27

OpenStudy (anonymous):

uve got three equations and three variables

OpenStudy (amistre64):

dont loan friends money ;)

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

help ur friendswith the maximum u can d0 :)